OpenStudy (sonyalee77):

Mr. Smith is thinking of a number. Three times the number minus 12 equals 36. What is the number Mr. Smith is thinking of? Identify /Explain the Guess, Check, and Revise Problem Solving Strategy. Also, list the seven steps you take while using this strategy.

OpenStudy (anonymous):

the answer is 16 dont know the steps though

OpenStudy (anonymous):

36+12=48 48/3= 16

OpenStudy (anonymous):

3x-12=36 3x=48 x=16